EXPERIMENTAL AND AB-INITIO STRUCTURE OF BrNO_{2}

\underline{F. KWABIA TCHANA}, J. ORPHAL, I. KLEINER, CNRS, Laboratoire de Photophysique Moléculaire, 91405 Orsay Cedex, France; H. D. RUDOLPH, Fachbereich Chemie, Universität Ulm, 89069 Ulm, Germany; H. WILLNER, P. GARCIA, Fachbereich C-Anorganische Chemie, Bergische Universität Wuppertal, 42097 Wuppertal, Germany; O. BOUBA, Faculté des Sciences, Université de Ngaoundéré, B.P. 454 Ngaoundéré, Cameroun; J. DEMAISON, Laboratoire de Physique des Lasers, Atomes et Molécules, UMR CNRS 8523, Université de Lille 1, 59655 Villeneuve d′Ascq Cedex, France; and B. REDLICH, FOM Institute for Plasma Physics, P.O.Box 1207, 3430 BE Nieuwegein, Netherlands.

The ν_{2} fundamental bands of different isotopomers of BrNO_{2} (^{79}Br^{15}N^{16}O_{2}, ^{81}Br^{15}N^{16}O_{2}, ^{79}Br^{14}N^{18}O_{2}, and 79Br^{14}N^{16}O^{18}O) located around 13 μm, were recorded using high-resolution Fourier-transform infrared spectroscopy. More than 8000 lines of all these isotopomers were reproduced using a Watson-type A-reduced Hamiltonian with a rootmean-square deviation of better than 7 × 10^{−4} cm^{−1}, for the four isotopomers. Rotational and centrifugal distortion constants for the v_{2} = 1 states as well as for the vibrational ground states of these isotopomers have been determined. For the first time, an analysis of the ground state rotational constants obtained in this study combined with the constants obtained in our previous work on the ν_{2} bands of ^{79}Br^{14}N^{16}O_{2} and ^{81}Br^{14}N^{16}O_{2}^{a}, allow us to calculate the r_{m}-structure of nitryl bromide. A new ab-initio structure of nitryl bromide calculated at the CCSD(T)/SDB-aug-cc-pVQZ level of theory is presented and found in fair agreement with the experimental structure.

aF. Kwabia Tchana, J. Orphal, I. Kleiner, B. Redlich, R. Mbiake and O. Bouba, J. Mol. Spectrosc., 216, 2, 292-296 (2002).